A basketball of mass 0.0400 kg is dropped from a height of 5.00 meters to the ground and bounces back to a height of 3.00 meters .


A. On its way down, how much potential energy does the ball lose?


B. On its ways, how much potential energy does the ball regain?

1
Expert's answer
2022-01-07T09:30:07-0500

A. The potential energy is defined as follows:


"E_1 = mgh_1"

where "m =0.0400kg" is the mass of the ball, "g = 9.81m\/s^2" is the gravitational acceleration, "h_1 = 5.00m" is the initial height. Thus, obtain:


"E_1 = 0.04\cdot 9.81\cdot 5 \approx 1.96J"

B. Similarly, on the way up:


"E_2 = mgh_2 = 0.04\cdot 9.81\cdot 3 \approx 1.18J"

Answer. A. 1.96J, B. 1.18J.